OpenStudy (anonymous):

The perimeter of a right triangle is 24 cm. How long are its sides? Medal goes to best answer :)

OpenStudy (anonymous):

not enough information

OpenStudy (anonymous):

its literally the problem for my math class :/

OpenStudy (anonymous):

should i just say "not enough information"?

OpenStudy (anonymous):

well it is true that \(6+8+10=24\) and that \(6,8,10\) are sides of a right triangle

OpenStudy (anonymous):

i guess you were just supposed to know this

OpenStudy (anonymous):

well it's isoceles so that means two legs are equal lengths, and there is a 90 degree angle

OpenStudy (anonymous):

and perimeter is length around so that means 2x + y = 24?

OpenStudy (anonymous):

it says nothing about isosceles

OpenStudy (anonymous):

it says "right" triangle

OpenStudy (anonymous):

oh oops, it does say isosceles in parentheses i'm sorry

OpenStudy (anonymous):

so should i just say that it has something to do with the equation 2x + y = 24

OpenStudy (anonymous):

ok can you post the exact problem? maybe there is enough information

OpenStudy (anonymous):

"The perimeter of a right isosceles triangle is 24 centimeters. How long are its sides?"

OpenStudy (anonymous):



OpenStudy (anonymous):

\[2x^2=y^2\] and \(2x+y=24\)

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

you good from there?